Technical Staff Blog

Archives May 2016

Last update on .

network-2400px.png 09:53: We are currently experiencing a widespread network outage with any site outside the CS department, including wireless. Tstaff is trying to confirm it's not an issue on our end and we have reached out to the CIS networking team to get their help debugging this.

10:06: We lost the network link ...

Last update on .

network-2400px.png On May 31st @ 7am CIS will be shutting down one of our two distribution switches and, because it is directly attached, we will be shutting down our backup firewall. Because of redundancy, we do not expect this work to cause any network outages. CIS will spend the remainder of May 31st moving the switch to ...

Last update on .

ibm-gpfs.jpg Overall, the migration of GPFS to the basement went remarkably well! We lost a total of two power supplies, both failures we could sustain through redundant hardware, and ran into a few bumps along the road, but it could have been a catastrophic failure and sure wasn't. We do have some additional cleanup that ...

Last update on .

ibm-gpfs.jpg The GPFS hardware is scheduled to be moved starting May 23rd @ 06:30am. Please check this page on the day of the move for periodic status updates...

... and so it begins...

05:45: Prep work for the move has begun. We are working to ensure that the most critical services remain active during the file ...

Last update on .

postfix.png As part of our ongoing virtualization efforts, we rolled out an upgraded version of our mail relay server yesterday at noon. Unbeknownst to us, the default value for a configuration variable changed in a subtle way that affected delivery to some subdomains. This definitely affected delivery to some of our email lists. We put a ...

Last update on .

ibm-gpfs.jpg In just under two weeks, Tstaff will be physically moving our GPFS file system hardware from CIT 531 to the datacenter in the basement of the CIT. The purge of the upstairs machine rooms is the culmination of nearly two years of work by Tstaff. We will be shutting down GPFS @ 6:30am to begin ...

Last update on .

We had an emergency failover of the CS firewall at 4:45pm because two security updates screwed up the active firewall software configurations.  The fail over was a manually process because some of active firewall software were removed due to a package conflict with the security updates.  There was about a ten minutes outage for ...

Last update on .

Mailing List Issues

Some CS mailing list messages were not immediately archived over the weekend of April 29 - May 1.  A filled disk caused archiving of the messages to be delayed.  As we fixed the problem this evening, we inadvertantly caused many of those message to be re-sent, so some users may notice that they received duplicate messages ...

Daily archives

Previous month

April 2016

Next month

June 2016

Archives